OpenStudy (anonymous):

1.The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (anonymous):

original SA = 2( wh + lw + lh) when all sides are tripled: new SA = 2 (3w*3h + 3l *3w + 3l*3h) = 2 ( 9wh + 9lw + 9lh) = 18 ( wh + lw + lh) comparing original SA with New SA 2( wh + lw + lh) :18 ( wh + lw + lh) 2:18 1:9 therefore original SA will increase 9times i.e 9*150 =1,350cm²